OpenStudy (anonymous):

Can someone please help me calculate this. the minimum safe distance between vehicles on a highway is the distance a vehicle can travel in 2.0 s at a constant speed. Assume that a 1.2 × 103-kg car is travelling 72 km/h [S] when the truck ahead crashes into a northbound truck and stops suddenly. I know that the acceleration is 10m/s^2 But then I don't know which formula to use.

OpenStudy (anonymous):

What do you want to know?

OpenStudy (anonymous):

I need to know if the car is at the required safe distance behind the truck, what is the separation distance? Would I just use the formula, d=vit+ 1/2a(t)^2?

OpenStudy (anonymous):

First of all. Where did you get that acceleration? Also, If you don't have the distance between the car and truck, you can't know if the car is in the required safe distance or not.

OpenStudy (anonymous):

To calculate the acceleration first I did a=vi-vf/t which is 0-(-20m/s)/2.0s=10m/s^2. I thought you just need to calculate the distance between the cars to actually know if they are at a safe distance?

OpenStudy (anonymous):

I think you're right. It doesn't give you the distance that the truck is at so you can't calculate the difference

OpenStudy (anonymous):

Huh. You're right about the acceleration. Although, the braking force required is kind of high. You can calculate the safe distance between the objects utilizing the equation: \[d = Vt\] Where V is the constant velocity of the car.

OpenStudy (anonymous):

Oh okay, Thank you.